I need an equation which models how old in years you will be when you are 3 times as old as the younger person. Ages 41 and 12
Answer by rwm(914)

(41+x)=3*(12+x)
x=2.5
In 2.5 years I will be 3 times as old as the younger person.
41+2.5=43.5
12+2.5=14.5
3*14.5=43.5